OpenStudy (anonymous):

Saturn is approximately 1.5 billion kilometers from the sun. What is this distance in AU? 0.1 AU 10 AU 100 AU 1,000 AU

OpenStudy (anonymous):

I answered a similar question here: http://openstudy.com/study#/updates/5013018fe4b0fa24672fdae0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Did you get an answer?

OpenStudy (anonymous):

nope :/

OpenStudy (anonymous):

1.5 x 10^9 km (1 AU)/(1.5 x 10^8 km) = ?

OpenStudy (anonymous):

The first number is the distance given in your problem: 1.5 billion

OpenStudy (anonymous):

\[(1.5 x 10^9)(\frac{ 1 AU}{1.5 x 10^8 km}) = ??\]

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

Did you get an answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Great!